slipknot3066:

Simplify 2(x + 4) + 3(x - 4) 5x - 4 5x 5x + 4

Gdeinward:

do you have any ideas?

slipknot3066:

no

Gdeinward:

First you must distribute each half of the expression

ZoeyBiocth16:

the right answer is 5x-4 because if you write the problem out like this: 2(x + 4) + 3(x - 4) 2(x+4) = 2x+8 2x+8 + 3(x-4) 3(x-4) = 3x-12 2x+8 +3x-12= 5x-4

Gdeinward:

I was trying to go one on one to help slipknot each step of the way

ZoeyBiocth16:

doing that takes longer tbh

Gdeinward:

It does but its worth it.

ZoeyBiocth16:

its better if you tell what th answer is then explain or show how you got it

ZoeyBiocth16:

the*

Gdeinward:

true. But there is more than one way to skin a cat

ZoeyBiocth16:

true

Gdeinward:

THey could disregard the solution and look at the answer. It helps to engage the person so they know what they are doing, but I can respect your method.

ZoeyBiocth16:

i mean im just going off of what i was told in the past but every method is acceptable except for just telling the answer and not explaining how you got it

slipknot3066:

thanks XD

Gdeinward:

I have been told the same. I just expanded on it a little.

ZoeyBiocth16:

your welcome slip

Gdeinward:

Your welcome!